OpenStudy (anonymous):

Find the domain of the function and express the answer in interval notation: f(x)=3x-9

OpenStudy (anonymous):

(infinity,-infinity) as the answer

OpenStudy (zarkon):

really?

OpenStudy (zarkon):

\[(-\infty,\infty)\]

OpenStudy (anonymous):

i really dont know how to do this, i just learn how to figure it out when there is a square root, and i can figure it when it is dividing. I havent had a math class since 2005. so its taken me some time to figure this out. im sorry

OpenStudy (zarkon):

in interval notation the smaller number goes to the left

OpenStudy (anonymous):

didnt you just write the same answer as me? i just used words

myininaya (myininaya):

no

myininaya (myininaya):

\[(-\infty,\infty)\]

OpenStudy (anonymous):

why does the smaller number go to the left? is it because the negative is on the left with the graph for x-values?

myininaya (myininaya):

we list the smaller number first because it comes before the bigger number

myininaya (myininaya):

you know its like writing a set when we write the natural numbers we start with the smaller one N={1,2,3,...} because we want to show a pattern (-inf,inf) is just a set of numbers as well

OpenStudy (anonymous):

sad part is, i know the smaller number is first then the bigger number, thats logic and i should of remember that. i just dont understand how to find the domain and write the answer in interval notation.

myininaya (myininaya):

anything in this form \[f(x)=a_nx^n+a_{n-1}x^{n-1}+a_{n-2}x^{n-2}+ \cdot \cdot \cdot +a_1x+a_0\] is called a polynomial and a polynomial is defined for all real numbers and therefore the domain is all real numbers using interval notation we say: \[(-\infty,\infty)\]

myininaya (myininaya):

all those a's are just constants by the way

myininaya (myininaya):

all those should be positive integer exponents

myininaya (myininaya):

or neutral integer exponents lol

OpenStudy (anonymous):

i know the form, my teacher showed it to because i told him i didnt understand this. i can figure this out when dividing and when there is a squareroot.

OpenStudy (zarkon):

n just needs to be a non-negative integer

myininaya (myininaya):

right

myininaya (myininaya):

positive or neutral

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

tell your teacher that all these problems are bull hockey

OpenStudy (anonymous):

a function is supposed to come with the domain as part of the definition.

OpenStudy (anonymous):

questions that say "find the domain of f" are so much non-sense. if it said "find the largest possible set on which f can act" then it might make some sense

OpenStudy (anonymous):

just thought i would mention it

OpenStudy (anonymous):

yeah i dont think im going to say that to my college teacher...

OpenStudy (zarkon):

if they are any good they would agree with you ;)

OpenStudy (anonymous):

i understand everything else that he throws at me, but this i dont get

OpenStudy (anonymous):

thanks for the help, i think the these 4 questions i asked about these with help me finish my homework.

OpenStudy (anonymous):

no i was kidding when i said "tell your teacher..." when you see a function and are asked for the domain, start by thinking "all real numbers"

OpenStudy (anonymous):

like "innocent until proven guilty"

myininaya (myininaya):

lol

OpenStudy (anonymous):

then you usually only have to worry about two things, at least until you get further 1) no zeros in the denominator 2) no negative numbers inside a radical

OpenStudy (anonymous):

ok thanks, i kinda figured that you were kidding, plus i only have to put with this class for 7 more day...lol

OpenStudy (anonymous):

7 more days then freedom!

OpenStudy (anonymous):

only fret about those things. if you see no radicals, no denominators, say "all real numbers" and have a nice day. that is it

OpenStudy (anonymous):

if you see a denominator, make sure it is NOT zero if you see a radical, make sure what is inside is greater than or equal to zero. finis

OpenStudy (anonymous):

you have a nice day too